Six different positive numbers(excluding zero) are added together, and their sum is 11. Four of the numbers are 4, 3, 2, and 1. Find the other two numbers.
Answer by checkley75(3666)

11=4+3=2+1+x
11=10+x
x=11-10
x=1 For 2 positive numbers to add up to 1 they must be fractions or decimals. Thus
y+z=1
1/4+3/4=1 or
.25=.75=1 or
1/5+4/5=1 or
.2+.8=1 or
2/5+3/5=1
.4+.6=1 etc.